OpenStudy (anonymous):

how many times greater is .10% than .02%?

OpenStudy (campbell_st):

try 0.1 divided by 0.02

OpenStudy (anonymous):

not it

OpenStudy (kropot72):

Because they are both percentages you can divide one percentage by the other to find the answer: \[\frac{0.10}{0.02}=?\]

OpenStudy (anonymous):

i got 322.59%

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

.10%-.02%/.02%

OpenStudy (anonymous):

0.10 = 1/10

OpenStudy (anonymous):

0.02 = 1/50

OpenStudy (anonymous):

(1/10)/(1/50)

OpenStudy (anonymous):

Can you divide fractions?

OpenStudy (anonymous):

criss cross multiplication

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its asking how many times greater

OpenStudy (anonymous):

not the rate of change?

OpenStudy (anonymous):

I am hinting at how to do it. I am not going to tell you the answer outright.

OpenStudy (anonymous):

Remember that decimals are fractions.

OpenStudy (anonymous):

0.10 = 1/10 0.02 = 2/100 = 1/50

OpenStudy (anonymous):

yeEAS (Borat accent)

OpenStudy (campbell_st):

here is an online calculator just do 0.1 divided by 0.02 and then press the equals key http://web2.0calc.com/

OpenStudy (anonymous):

\[\frac{ 0.10 }{ 0.02} = \frac{\frac{ 1 }{10} }{ \frac{ 1 }{ 50 }}\]

OpenStudy (anonymous):

wait i hae to move the decmials if its percentage

OpenStudy (anonymous):

have

OpenStudy (anonymous):

Ohh I didn't see the percent.

OpenStudy (anonymous):

Remember to convert a percent to a decimal divide by 100.

OpenStudy (campbell_st):

? x 2 = 10 whats the value of ? thats all the question is asking.

OpenStudy (anonymous):

.10% = .001 .02%=0.0002

OpenStudy (anonymous):

hmmmm ok so i divide .001 by .0002?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

wait i thought i have to do something different

OpenStudy (anonymous):

Women in the United States continue to be paid less than men overall. Nationally, in 2007, the median weekly earnings of female full-time wage and salary workers totaled $614, while it was $766 for males. In 2007, male full-time wage and salary workers earned _________ % more than female full-time wage and salary workers.

OpenStudy (anonymous):

isnt the former problem similar to the this one

OpenStudy (anonymous):

this*

OpenStudy (anonymous):

if not whats the difference

OpenStudy (anonymous):

It's not a % value so you don't have to divide by 100.

OpenStudy (anonymous):

hmmmm maybe its cuz this is a absolute quantity and the former was a relative quantity?

OpenStudy (anonymous):

so in this case you would (766 – 614)/614 = 24.8%

OpenStudy (anonymous):

Well... I guess that's one way to put it....

OpenStudy (anonymous):

Yeah that's right.

OpenStudy (anonymous):

friggin confusing hella

OpenStudy (anonymous):

Haha. You will get the hang of it soon.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

wait what was the answer though LMAO it was 5?